why do we subtract the profit from loss? take example:
A Cement company earns a profit of RS 8 on selling  1 white cement bag and incurs a loss of RS 5 on selling 1 grey cement bag. the company sells 3000 bags of white cement and 5000 bags of gray cement in a month. WHAT IS PROFIT OR LOSS?
my question is that after finding loss and profit why do we subtract them.

Grade:7

1 Answers

Saurabh Koranglekar
3 years ago
Profit for selling each white cement bag = Rs. 8
Loss for selling each grey cement bag = Rs. 5

Bags of white cement sold by the company = 3,000
Bags of grey cement sold by the company = 5,000

Hence profit earned by the company = 3,000 × 8 = Rs.24,000
Loss incurred by the company = 5,000 × 5 = Rs. 25,000

25,000 is greater than 24,000, Hence there is a loss of Rs. 1000 for the company.